Some measurements to re-estimate the sensing noise of the PCals sensors, on both NE and WE (times given in UTC):
- 9h36: stop control loop and set laser power to 0
- 9h37: stop laser pump diodes --> sensing noise data from GPS 1408786680, for about 90 s.
- 9h39m35s: switch ON the LED with IRIG-B signal. Note: LED amplitude not stable on WE, while stable on NE)
- 9h41m35s: switch OFF the LED --> sensing noise data again.
- 9h43m00s: laser pump diode one
- 9h45m30s: laser power set to ~1.3 W --> data for free laser power noise
- 9h51m00: control loop closed.
At 9h53 UTC, started the script for PCAl sensor calibration: python3 /virgoDev/Automation/scripts/CALI_O4/pcal_sensing/send_commands.py 120. which worked well. Note: there was an error in one useless import from the script (#import pandas as pd), so this data have not been taken since a few months. From now, it will be again taken regularly when launching the calibration procedure with ITF unlocked.
From ~10h01 UTC, Pcal working in their nominal conditions.
- Figure 1 shows the nominal signals, with 1.3 W, closed loop on Tx_PD1.
- Figure 2 shows the sensins noise: same signals but with laser off.
- Figure 3 shows also the sensing noise but all superposed.
For WE, the Rx_PD1 sensing noise is around 1e-7 W/sqrt(Hz) at 100 Hz, but some additional noise at the 1e-6 W/sqrt(Hz) level is present when the laser is on.
For NE, the Rx_PD1 sensing noise is much larger, around 1e-6 W/sqrt(Hz) at 100 Hz.